OpenStudy (anonymous):

An inner copper ball (radius=2.0m) has a positive charge of6.00*10^-6C and the outer copper shell (inner radius=4.0m) has anequal but opposite charge. (a) Find the capacitance of thisarrangement. (b) find the energy stored in this capacitor.

OpenStudy (anonymous):



OpenStudy (samgrace):

\[V = Ed = \sigma/\epsilon_{0} \times d = Qd/\epsilon_{0}A\] \[C = Q/V \] and \[U = 1/2 Q^{2}/C \]